OpenStudy (anonymous):

an english instructor asserted that students test grades are directly proportional to the amount of time spent studying. Lisa studies for 9 hr for a particular test and gets a score of 75. at this rate, how many hours would she have had to study to get a score of 92? please help

OpenStudy (lgbasallote):

directly proportional so \[y = kx\] let x be the hours she take studying and y is her grade \[75 = 9k\] \[k = \frac{75}{9}\] now to get a score of 92... \[y = kx\] \[92 = \frac{75}{9} x\] solve for x got it?

OpenStudy (anonymous):

in the book the answer is 11.04 when i divide that i dont get that answer

OpenStudy (lgbasallote):

really i got 11.04

OpenStudy (lgbasallote):

\[92 = \frac{75}{9} x\] \[\frac{92 \times 9}{75} = x\]